OpenStudy (anonymous):

Hi there I have a ? I'm stuck on! Solve the following equation for all solutions 3x^-5=9375 answers accurate to 3 decimal places! (lesser)solution = (greater) solution =

OpenStudy (agent0smith):

this?\[\Large 3 x^{-5} = 9375\]have you learned much about complex numbers...?

OpenStudy (anonymous):

yea that's the equation! no not really! its new to me!

OpenStudy (agent0smith):

If you haven't learned imaginary numbers, then you'll only be able to find one solution. Bring the negative exponent in the denominator to make it positive \[\Large \frac{ 3 }{x^5 } = 9375\]can you try solving for x?

OpenStudy (anonymous):

would you add 1/-5 to each side?

OpenStudy (agent0smith):

No... Multiply both sides by x^5.

OpenStudy (anonymous):

meaning raising both sides

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

then do I root those?

OpenStudy (agent0smith):

I don't know what you did to get those... we're not putting both sides to the power of anything. Multiply both sides by x^5 \[\Large \frac{ 3 }{x^5 } *x^5= 9375x^5\] becomes \[\Large 3 = 9375 x^5\]

OpenStudy (agent0smith):

Then divide both sides by 9375.

OpenStudy (anonymous):

It doesn't make any sense to me when i divide both sides by 9375 i'm lost

OpenStudy (anonymous):

i thought I would have to divide by 3

OpenStudy (agent0smith):

You need to get x^5 alone. Dividing by 3 won't do that.

OpenStudy (anonymous):

so cancel out 3=x^5

OpenStudy (anonymous):

i get 1/3125

OpenStudy (agent0smith):

1/3125 = x^5 now you have to take the 5th root of both sides (ie take the fifth root of 1/3125 and that'll be your answer)

OpenStudy (anonymous):

which is 3.355 to 3 place decimal

OpenStudy (anonymous):

but how do I figure out lesser solution and greater solution

OpenStudy (agent0smith):

There's only one real solution. The others are all complex numbers...

OpenStudy (agent0smith):

And the fifth root of 1/3125 is 1/5.

OpenStudy (anonymous):

which is 3.355 right

OpenStudy (anonymous):

to the 3 place decimal

OpenStudy (agent0smith):

No... it's 0.2.

OpenStudy (anonymous):

oh ok i see it now!

OpenStudy (anonymous):

makes perfect sense now! thx so much I appreciate it